OpenStudy (anonymous):

NECESITO DERIVAR ESTO V(r)=k(20r^2-r^3); 0

OpenStudy (mathmale):

I suggest that you go to translate.google.com and obtain a translation of your question into English. Most discussion on OpenStudy is in English. With translate.google.com, it's really easy to switch back and forth between languages. Te sugiero que vayas a translate.google.com y obtener una traducción de su pregunta en Inglés. La mayor discusión sobre OpenStudy está en Inglés. Con translate.google.com, es muy fácil cambiar de un lado a otro entre las lenguas.

OpenStudy (mathmale):

You need the derivative of V(r)=k(20r^2-r^3). Necesita la derivada de V (r) = k (20r ^ 2-r ^ 3)

OpenStudy (mathmale):

\[\frac{ d }{ dr }V(r)=k\frac{ d }{ dr }(20r^2-r^3)\]

OpenStudy (mathmale):

y que es igual a\[20k\frac{ d }{ dr }r^2 -k \frac{ d }{ dr }r^3\]

OpenStudy (mathmale):

¿Puedes ahora terminar la tarea de encontrar la derivada?

OpenStudy (anonymous):

creo que si muchas gracias